OpenStudy (anonymous):

Fool's problem of the day, Today a very easy trigonometry problem, \( (1) \) If \( \cos ^{-1} x - \cos ^{-1} \frac y 2 = \alpha \), then find the value of \( 4x^2-4xy\cos \alpha +y^2\) PS:The answer is (probably) Google-able, so post a solution instead ;)

OpenStudy (anonymous):

When you all get done here please come help me with my question. Nobody seems to be helping anymore.

OpenStudy (anonymous):

\[\cos (\cos^{-1} x-\cos^{-1} y \div2)=\cos \alpha\] \[x \times y \div 2 + {(\pi-2x)\times(\pi-y)}\div4 =\cos \alpha\]

OpenStudy (anonymous):

its easy cos(inv) [x*y/2 + sqrt((1-x2)(1-y2/4))]=a sqrt((1-x2)(1-y2/4))]=cos a - x*y/2 ,now square both sides (1-x2)(1-y2/4)=cos^2 a + ((xy)^2)/4 - xycosa 1-y2/4 -x2 +(xy)^2/4=cos^2 a + ((xy)^2)/4 - xycosa 1-y2/4 -x2=cos^2 a - xycosa 4-y2-4x2=4cos ^2 a - 4xycosa 4x2-4xycosa+y2=4(1-cos^2 a)=4sin^2 a Q.E.D.

OpenStudy (anonymous):

is it okay @experimentX @FoolForMath

OpenStudy (experimentx):

hahah ... lol you seriously overrate me.

OpenStudy (anonymous):

"overrate"-what's that supposed to mean???

OpenStudy (experimentx):

like i would know the answer ... how did you get the first step??

OpenStudy (anonymous):

i tried many ways on paper but this method stood out

OpenStudy (anonymous):

Congratz @Sarkar You did it! :D

OpenStudy (experimentx):

I didn't understand the first step??

OpenStudy (anonymous):

I am thinking to make a Hall of Fame for the problem of the day, What say you guys?

OpenStudy (apoorvk):

yeah yeah. sure!

OpenStudy (experimentx):

I understand ... now.

OpenStudy (anonymous):

@experimentX - i just used the formula of cos(inv) x - cos(inv) y= cos(inv)[xy-sqrt(1-x2)(1-y2)]

OpenStudy (anonymous):

cos(inv) x - cos(inv) y= cos(inv)[xy+sqrt(1-x2)(1-y2)]

OpenStudy (anonymous):

Yes, that's Inverse circular function 101!

OpenStudy (experimentx):

wow ... there was a formula like that??? Congratz @Sarkar Taking Cos(...) = cos(a) yields the same.

OpenStudy (experimentx):

I took cos ... got your second step ... and stuck .. lol

OpenStudy (anonymous):

Yes @experimentX It's easy to prove too, Consider \( cos^{-1} x = A\) and \(\cos^{-1} y = B\) and then use \( \cos (A+B) \)

OpenStudy (anonymous):

Stuck why???@X

OpenStudy (experimentx):

hahaha ... I'll remember that. and also persistence is the key to success.

OpenStudy (anonymous):

Inverse circular function doesn't have much significance higher mathematics (I don't know why!) so it's somewhat least practiced.

OpenStudy (experimentx):

Well, I never learned that on school ... ;D

OpenStudy (anonymous):

i do agree @FFM but its in our course so have to brush it constantly...

OpenStudy (anonymous):

It's standard high school material here.

OpenStudy (anonymous):

@Sarkar JEE aspirant?

OpenStudy (anonymous):

yes@FFM

OpenStudy (anonymous):

and you???

OpenStudy (anonymous):

2013?

OpenStudy (anonymous):

yup...

OpenStudy (anonymous):

I failed long back :P

OpenStudy (anonymous):

Good luck! :)

OpenStudy (experimentx):

Good Luck!!

OpenStudy (anonymous):

are you an Indian.???

OpenStudy (anonymous):

@FFM

OpenStudy (anonymous):

thanks for your wishes

OpenStudy (anonymous):

Yes of-course :)

OpenStudy (anonymous):

didnt know that until now,,lol

OpenStudy (anonymous):

lol, I don't usually publicize my nationality unless asked to do so ;)

OpenStudy (anonymous):

whatever,can you help me with a sum???

OpenStudy (apoorvk):

Foolformath can help you with ANY sum!! :)

OpenStudy (anonymous):

Okay guys, hope you all enjoyed today's problem. See you all very soon. @Sarkar Sorry, I have to leave now, you can you post it via "Ask the question..." feature. I will check it later :) Just don't forget to type @FoolForMath. Since, @FFM don't notify me.

OpenStudy (apoorvk):

ohh.. ffm is some other username. i see. lol

OpenStudy (experimentx):

be back with Hall of Fame problem!!!

OpenStudy (anonymous):

it goes something like this \[\sin^{-1} [(a+bcosx)/(b+acosx)]\]...diffrentiate it i know how to diff it but is there a possible sub we can use here????

OpenStudy (anonymous):

will do @FoolForMath

OpenStudy (experimentx):

@Sarkar if you put that into b/a - ((b^/a)-b)/(b+acosx) .... it might make simpler.

OpenStudy (experimentx):

* b/a - ((b^2/a)-b)/(b+acosx)